OpenStudy (anonymous):

Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work.

OpenStudy (anonymous):

If they the measurements of any right triangle, then Pythagoras's Theorem must apply. The square of the longest side is equal to the sum of the squares of the other two side. \[c^2 = a^2 + b^2 \] \[51^2 = 24^2 + 45^2 \]

OpenStudy (anonymous):

If they are equal, the both sides, then the side measurements are of a special right triangle.

OpenStudy (anonymous):

thanks